# Parallelized Planning-Acting for Efficient LLM-based Multi-Agent Systems

### v1

@@ 3.1 Experiment Setup
All experiments are conducted in the Minecraft Java Edition version 1.19.4 gaming environment using the Qwen-Plus model  Qwen et al. (2024), while multi-modal experiments utilized the Qwen-VL-Plus model  Bai et al. (2023). The game server operates continuously without pausing during interactions with LLMs, thereby necessitating that all agents perform in real-time. All evaluation metrics are captured dynamically during these interactions. Metrics such as ’time’ denote the total duration required by the multi-agent systems to complete their designated tasks, encompassing the time expended on communications with LLMs.
3.2 Benchmark Task Scenarios
To evaluate the generalizability of our multi-agent framework in Minecraft, we designed three distinct task scenarios: Resource Collection, Boss Combat, and Adversrial Player versus Player (PvP). These scenarios were selected to cover a range of application domains, showcasing the versatility of our multi-agent framework across diverse contexts.
  • Resource Collection: This task involves gathering resources in the Minecraft world, evaluating the efficiency of the comprehensive skill library and the coordination of multi-agent system in performing repetitive and structured tasks.
  • Boss Combat: This task requires the multi-agent system to defeat a strong boss-level enemy in a predefined, highly intricate combat environment, evaluating advanced strategic planning, real-time decision-making, effective collaboration among agents and the system’s ability to handle dynamic and challenging environments.
  • Adversarial PVP: This task focuses on team-based adversarial competition, where two teams of agents compete against each other in a combat scenario. The primary objective is to evaluate the agents’ ability to strategize, adapt, and outperform opponents.
3.3 Main Results
3.3.1 Resource Collection Task
We first evaluate the effectiveness of our skill library’s recursive task decomposition mechanism by assessing a single agent’s performance on basic resource collection tasks in Minecraft. This initial evaluation aimed to verify the comprehensive skill library’s capability in automating complex workflows, thereby validating its efficiency and reliability in handling fundamental Minecraft tasks. As shown in Table 1, when the recursive task decomposition mechanism is ablated, the system can only complete short-term tasks that require fewer steps and shows a reduction in efficiency. In contrast, utilizing our comprehensive skill library enables efficient completion of all tasks.
Table 1: Completion time and success rate (SR) comparison with and without the Recursive Task Decomposition Mechanism (RTDM). All results are shown with the mean and standard deviation over 10 trials.
Task with RTDM w/o RTDM
Time(min) SR Time(min) SR
[Uncaptioned image] $0.3\pm 0.2$ $100\%$ $2.8\pm 2.5$ $100\%$
[Uncaptioned image] $0.6\pm 0.3$ $100\%$ $4.6\pm 2.5$ $80\%$
[Uncaptioned image] $1.4\pm 0.5$ $100\%$ $N/A$ $0\%$
[Uncaptioned image] $4.7\pm 1.3$ $100\%$ $N/A$ $0\%$
[Uncaptioned image] $6.2\pm 1.6$ $100\%$ $N/A$ $0\%$
We then evaluate our framework’s performance on eight more complicated resource collection tasks. Table 2 compares the completion times between our multi-agent system (3 agents) and single-agent baseline, demonstrating the efficiency gains through coordination.
@@ 3.3.1 Resource Collection Task
Table 2: Completion time (in minutes) of resource collection tasks comparison between multi-agent(MA) and single-agent(SA) systems. All results are shown with the mean and standard deviation over 10 trials. For detailed task definitions, please refer to Appendix C.
Task MA Time SA Time
@@
The experimental results demonstrate that our Minecraft comprehensive skill library, when employed within a multi-agent framework, can efficiently complete various resource collection tasks as well. As shown in the comparisons, the system with three agents significantly reduces task completion times compared to the single-agent baseline across most tasks, validating the effectiveness of our approach and showcasing the efficiency benefits of multi-agent collaboration. However, the observed non-linear scaling (multi-agent time > single-agent time/3) does not always achieve ideal linear speedup due to two inherent constraints:
  • Resource dependency chains: Sequential prerequisites in crafting workflows (e.g., iron tools required for diamond mining)
  • Spatial contention: Overlapping access to resources in close proximity (e.g., multiple agents competing to gather items from the same ore cluster)
3.3.2 Boss Combat Task
We conduct comprehensive evaluations across three challenging combat scenarios with varying agent team sizes. Experiments are conducted in three predefined scenarios, each featuring an extremely powerful and representative boss from Minecraft’s major dimensions: the Elder Guardian (Overworld), the Wither (Nether), and the Ender Dragon (End). The initial inventory of agents remained consistent across all scenarios. Their equipment are validated to be collectible within a short time during the resource collection task, and other consumable resources can also be efficiently gathered ultilizing our skill library. For further details, please refer to Appendix C.
@@ 3.3.2 Boss Combat Task
Table 3: Boss combat performance across three task scenarios. All results are shown with the mean and standard deviation over 12 trials. For detailed evaluation metrics definitions, please refer to Appendix C.
Task Scenario # Agents Time(min) Health Ratio Progress SR
@@
3.3.3 Adversarial PVP Task
In the adversarial player-vs-player (PVP) task, we conduct a direct comparison between the parallelized and serialized frameworks across various team sizes. As shown in Fig. 3, this setup clearly demonstrate and quantify the differences of these two methodologies under competitive conditions.
As shown in Fig. 3, the parallelized framework demonstrates a significant advantage over the serialized framework in this dynamic adversarial scenario. Our analysis indicates that this advantage is primarily attributed to our interruption mechanism, which enables agents to dynamically adjust their strategies and respond promptly to changes in the environment (e.g., seamlessly switch attack targets, prioritize health restoration, etc.).
Figure 3: Comparison of parallelized vs. serialized frameworks in PVP tasks. All results are shown with the mean and standard deviation over 10 trials.
3.4 Ablation Study
@@
  • w/o Parallelized Planning-Acting Framework: Replaced our concurrent architecture with traditional serialized execution (LLM call → action execution), disabling action interruption capabilities once execution begins.
@@
  • w/o Centralized Memory System: Disabled real-time team observation polling, team chat logs, action history and global progress information, thereby restricting agents to rely solely on their individual observations.
  • Multimodal Observation: Replaced text-based observations with visual inputs using Mineflayer’s prismarine-viewer, capturing first-person screenshots processed by a vision-language model (VLM) as observation.
These ablation configurations systematically evaluate the contribution of each architectural component to overall system performance, providing insights into their relative importance in complex multi-agent coordination tasks.
Experiment results shown in Fig. 4 highlight the critical roles of the parallelized planning-acting framework and centralized memory system in maintaining this high performance. When we replaced the text-based observation space with a multimodal observation system, we observed that although the inclusion of visual language models (VLMs) increased response latency and lowered the observation accuracy, the agents still performed reasonably well. This indicates that our framework remains robust even when handling different modalities.
@@
Figure 4: Ablation study in boss combat tasks across 12 trials in three scenarios: Elder Guardian (EG), Wither (Wi), and Ender Dragon (ED), where efficiency is defined as the inverse of completion time in minutes.
3.5 Human-Agent Interaction
Our framework also supports flexible human-agent interaction, allowing human players to either actively participate in task execution alongside agents, facilitating dynamic cooperation within the game environment, or to take on a guiding role by providing instructions to agents for task division, coordination, and strategic decision-making. Further experimental cases demonstrating the effectiveness of human-agent interaction are presented in Appendix C.4.
4 Related Works
@@
LLM-based Minecraft Agents. The development of LLM-based AI agents in Minecraft has evolved through several key approaches: Voyager  Wang et al. (2023) established the first LLM-based agent with automatic skill discovery using GPT-4  Achiam et al. (2023). Subsequent studies enhanced agents via specialized memory mechanisms  Park et al. (2024); Li et al. (2024a), specialized LLM fine-tuning  Feng et al. (2023); Zhao et al. (2024); Liu et al. (2024), task decomposition and causal graph learning  Yuan et al. (2023); Zhu et al. (2023); Yu and Lu (2024), and combination with reinforcement learning  Li et al. (2024b, 2023). Additionally, multi-modal information perception and processing were explored  Zheng et al. (2023); Cai et al. (2024), along with other novel techniques  Wang et al. (2024); Zhou et al. (2024). The development of benchmarks for general capabilities progressed with MineDojo  Fan et al. (2022) and MCU  Lin et al. (2023), while specific agent capabilities were assessed through additional benchmarks  Qin et al. (2024); Dong et al. (2024).
@@
LLM-based Multi-Agent Systems. Recent surveys  Mou et al. (2024); Guo et al. (2024) provide a comprehensive overview of the key advancements in LLM-based multi-agent systems. Building on these foundations, recent research has focused on several core areas: infrastructure frameworks for efficient agent coordination  Gong et al. (2024); Chen et al. (2024b); Zhang et al. (2024b); Hong et al. (2023); Zhang et al. (2024a), which introduce novel paradigms for task management and team collaboration; benchmark development to evaluate multi-agent performance in dynamic environments  Chen et al. (2024a); Dong et al. (2024), which has created robust testing environments to assess the generalization and efficiency of LLM-powered agents; large-scale social simulations  AL et al. (2024); Park et al. (2023); Yang et al. (2024), which explore how multi-agent systems model complex societal behaviors; and domain-specific applications  Wu et al. (2024); D’Arcy et al. (2024); Chen et al. (2024c) demonstrating the effectiveness of LLM-based agents in specific scenario simulation.
5 Conclusion
@@
We propose a novel parallelized planning-acting multi-agent framework that significantly enhances the responsiveness and adaptability of LLM-based MAS in dynamic environments. Our framework’s dual-thread architecture with interruptible execution mechanism enables real-time interaction and continuous adaptation, overcoming the limitations of traditional serialized execution paradigms. The comprehensive skill library and automated task decomposition mechanism further improve efficiency and coordination as an engineering contribution. Through extensive experiments, our framework demonstrates superior performance in both collaborative and adversarial scenarios, and achieves more immediate human-agent interaction, highlighting its potential for broader applications.
6 Limitations
While demonstrating significant advantages, our framework has three main limitations. First, the continuous LLM request design leads to higher computational costs, which may limit scalability in resource-constrained environments. Second, LLM hallucinations occasionally result in irrational action interruptions, affecting the reliability of agent behavior. Third, although our framework adapts to multimodal observations, the current fusion method relies solely on visual-language models (VLMs), potentially failing to fully exploit the complementarity of multimodal information and limiting performance in complex scenarios. For future work, we aim to optimize computational efficiency by refining LLM request mechanisms, mitigate hallucination effects through enhanced verification strategies, and improve multimodal fusion techniques to better leverage diverse sensory inputs.
@@ Appendix A Centralized Memory System
The memory of the multi-agent system is stored in JSON files, and agents read and modify their contents by reading and writing to these files. Below is an example of how the memory is stored:
A.1 Agent Observation & Status
"time": "day",
"health": 12.8,
"hunger": 17,
"position": "(10, 65, -8)",
"equipment": [
null,
null,
null,
null,
"bow",
null
"inventory": {
"arrow": 187,
"golden_apple": 4,
"diamond_pickaxe": 1,
"diamond_sword": 1,
"diamond_helmet": 1,
"diamond_chestplate": 1,
"diamond_leggings": 1,
"diamond_boots": 1,
"bow": 1
"nearbyBlocks": [
"end_stone",
"bedrock"
"nearbyEntities": "enderman, ender_dragon",
"nearbyPlayers": [
"Steve",
"Brine",
"Jeb",
"Notch"
A.2 Team Chat Log
"Day 1 - Morning": [
"SystemInfo: Number of End Crystals Remaining: 10",
"SystemInfo: Boss Health: 200",
"SystemInfo: Boss Health: 200",
"Notch: Lets prioritize taking down the end_crystals to prevent the Ender Dragon from healing. I will start attacking the nearest end_crystal with my bow. Everyone, cover me and be ready to assist if needed.",
"Steve: Im ready with my bow and arrows. Lets focus on taking down the end_crystals first to prevent the Ender Dragon from healing. Ill start targeting the nearest end_crystal. Stay alert for any incoming endermen but avoid engaging them unless necessary.",
"SystemInfo: Boss Health: 200",
@@ A.2 Team Chat Log
"Brine: Brine: Lets prioritize destroying the end_crystals to prevent the Ender Dragon from healing. Notch and I will start targeting the nearest crystal. Steve, Alex, and Jeb, please spread out and assist in taking down other crystals. Stay alert for any Endermen but avoid engaging them unless necessary. Report if you need help or spot any crystals.",
"Jeb: Im ready with my bow and arrows. Lets prioritize the end_crystal on the obsidian pillars to prevent the Ender Dragon from healing. Ill start targeting the nearest crystal. Stay alert for any incoming endermen but avoid engaging unless necessary."
A.3 Team Observation Log
"obs_entries": [
"Notch: Team, Im fully equipped with diamond armor, a bow, and 200 arrows. I see an end_crystal and an ender_dragon nearby, along with some endermen. Brine, Jeb, Steve, and Alex are also close. Lets focus on taking down the end_crystal first to prevent the Ender Dragon from healing. Everyone, prepare your bows and stay alert for any incoming endermen. Ready when you are!",
@@ A.3 Team Observation Log
"Brine: Team, Im currently well-equipped with a bow, 200 arrows, and diamond armor. The ender_dragon is nearby but hasnt attacked yet. I see Steve, Notch, Alex, and Jeb around. Lets focus on taking out the end_crystals first. I suggest we spread out to cover more ground and target the crystals from a distance. Stay alert for any Endermen, but avoid engaging them unless necessary. Report if you spot any crystals or need assistance.",
"Jeb: All end_crystals have been destroyed. Focusing on the Ender Dragon now. Everyone, keep your distance and use bows to attack from a safe range. Stay alert for any incoming endermen but prioritize the Ender Dragon. Lets finish this!",
"Steve: Im in critical condition and need immediate assistance! My health is down to 0. Brine, Notch, Jeb, please cover me or help me get to safety. I have a full bow and golden apples but cant fight effectively in this state. Endermen are nearby, be cautious!",
A.4 Team Action Log
"Notch": [
@@ A.4 Team Action Log
Part of above information is provided as human prompts for agents to plan actions. In addition to this, agents are also provided with identity information such as their name, team affiliation, and teammates’ names, to simulate a human-like multi-agent collaborative scenario.
Appendix B Minecraft Comprehensive Skill Library

### v2

@@ 3.2. Experimental Setup
All experiments are conducted in the Minecraft gaming environment using the Qwen-Plus model Qwen et al. (2024), while multi-modal experiments utilized the Qwen-VL-Plus model Bai et al. (2023). The game server operates continuously without pausing during interactions with LLMs, so all agents perform in real time. See Appendix for more details.
$3.$ Main Results
3.3.1. Resource Collection Task
We evaluate our framework’s performance on eight composite resource collection tasks in Minecraft. Each task requires agents to collect multiple items, where each item involves a long chain of prerequisite dependencies. For instance, crafting Diamond Armor necessitates first gathering wood, stone, and iron to produce the required tools, and all of which can be automatically accomplished through the recursive decomposition of our skill library. Agents only need to collaborate at the skill level. Table 1 compares the completion times between our multi-agent system and the single-agent baseline, demonstrating the efficiency gains achieved through coordination.
@@ 3.3.1. Resource Collection Task
Table 1. Average completion time and standard deviation of resource collection tasks comparing multi-agent (MA, fixed at 3 agents) and single-agent (SA) systems over 10 trials.
Task MA Time (min) SA Time (min)
@@
The experimental results demonstrate that our Minecraft comprehensive skill library, when employed within a multi-agent framework, can efficiently complete various resource collection tasks. As shown in the comparisons, the system with three agents significantly reduces task completion times compared to the single-agent baseline across most tasks, validating the effectiveness of our approach and showcasing the efficiency benefits of multi-agent collaboration. However, the performance gain does not scale linearly with the number of agents, which mainly stems from sequential dependency chains and spatial contention.
3.3.2. Boss Combat Task
We conduct comprehensive evaluations across three challenging combat scenarios, each involving different agent team sizes and featuring an extremely powerful and representative boss in Minecraft: Elder Guardian, Wither, and Ender Dragon. Each task requires agents to dynamically adjust their strategies based on environmental changes. For instance, when fighting the Wither, agents must monitor whether it has entered its final phase where it becomes immune to ranged attacks; when battling the Ender Dragon, agents need to first destroy the end crystals that continuously restore its health. Prior to each combat scenario, we equip each agent with identical initial supplies, including weapons, armor, and consumables. Our proposed comprehensive skill library supports the collection and crafting of these supplies, for example, the crafting of Diamond Armor has already been validated in previous experiments. We also observe that although each agent is homogeneous at initialization, heterogeneous division of labor can emerge through multiple rounds of actions and communication within the multi-agent system.
@@ 3.3.2. Boss Combat Task
Table 2. Performance of boss combat across three task scenarios, reporting mean values and standard deviations of multiple evaluation metrics over 12 trials. ”#Agents” refers to the number of agents. ”Time” refers to the minutes taken to complete the combat and achieve victory and is calculated only for successful trials. ”Health Ratio” refers to the ratio of the remaining health value of the team to its full health. ’Progress’ refers to the percentage of damage dealt to the boss to its full health. See Appendix for detailed metric definitions.
Task Scenario # Agents Time (min) Health Ratio Progress Success Rate
@@
20 $2.9\pm 0.8$ $67.9\pm 13.6\%$ $\pm 0.0\% 100.0100.0\%$
3.3.3. Adversarial PVP Task
Similar to the Boss Combat Task, in Adversarial PVP Task, agents were provided with initial combat resources, and then the two teams of agents engaged in battles with each other. We conduct a direct comparison experiment between the parallelized and serialized frameworks across various team sizes. Results are recorded when one team of agents is entirely defeated. After multiple experiments, victory rates and other relevant metrics are calculated. To promote fair confrontation, we provide each team with exactly the same initial supplies. As shown in Fig. 3, this setup clearly demonstrate and quantify the differences of these two methodologies under competitive conditions. The parallelized framework demonstrates a significant advantage over the serialized framework in this dynamic adversarial scenario. Our analysis indicates that this advantage is primarily attributed to our interruption mechanism, which enables agents to dynamically adjust their strategies and respond promptly to changes in the environment (e.g., seamlessly switch attack targets, prioritize health restoration, more efficient communication, etc.).
Figure 3. Comparison of parallelized vs. serialized frameworks in PVP tasks. All results are shown with the mean and standard deviation over 10 trials.
3.4. Ablation Study
3.4.1. Recursive Task Decomposition
We first evaluate the effectiveness of our skill library’s recursive task decomposition mechanism by assessing a single agent’s performance on resource collection tasks in Minecraft. This evaluation aimed to verify the comprehensive skill library’s capability in automating complex workflows, thereby validating its efficiency and reliability in handling fundamental Minecraft tasks. As shown in Table. 3, when the recursive task decomposition mechanism is ablated, the system can only complete short-term tasks that require fewer steps and shows a reduction in efficiency. In contrast, utilizing our comprehensive skill library with the recursive task decomposition mechanism enables efficient completion of all tasks.
Table 3. Average completion time with standard deviation and success rate (SR) comparison with and without the recursive task decomposition mechanism (RTDM). Tasks from top to bottom: (1) crafting table, (2) wooden tool, (3) stone tool, (4) iron tool, (5) diamond.
Task with RTDM w/o RTDM
Time (min) SR Time (min) SR
[Uncaptioned image] $0.3\pm 0.2$ $100\%$ $2.8\pm 2.5$ $100\%$
[Uncaptioned image] $0.6\pm 0.3$ $100\%$ $4.6\pm 2.5$ $80\%$
[Uncaptioned image] $1.4\pm 0.5$ $100\%$ $N/A$ $0\%$
[Uncaptioned image] $4.7\pm 1.3$ $100\%$ $N/A$ $0\%$
[Uncaptioned image] $6.2\pm 1.6$ $100\%$ $N/A$ $0\%$
In the more complex resource collection task we propose in Section 3.3.1, the MAS will not be able to make any meaningful progress without the recursive task decomposition mechanism, suggesting that invoking an LLM to independently plan the task dependencies embedded within a DAG would not be as efficient as utilizing a fixed mechanism.
3.4.2. Parallelized Framework
@@
  • w/o Parallelized Planning-Acting Framework (PPA): Replaced our parallelized framework with traditional serialized execution, disabling the interruption mechanism.
@@
  • w/o Centralized Memory System (CMS): Disabled real-time team observation polling, chat logs, action history and global progress information, thereby restricting agents to rely solely on their individual observations.
Experiment results shown in Fig. 4 highlight the critical roles of the parallelized planning-acting framework and the centralized memory system. Besides these ablation studies, our method also performs exceptionally well in direct comparative experiments. Please refer to Appendix for details.
@@
Figure 4. Ablation study in boss combat tasks across 12 trials in three scenarios: Elder Guardian (EG), Wither (Wi), and Ender Dragon (ED), where efficiency is defined as the inverse of completion time in minutes.
Table 4. Performance of boss combat across three task scenarios using different observation modalities, with a fixed team size of 10 agents. Reporting mean values and standard deviations of multiple evaluation metrics over 12 trials.
Task Scenario Observation Modality Time(min) Health Ratio Progress Success Rate
Elder Guardian Text $1.2\pm 0.3$ $86.9\pm 8.4\%$ $\pm 0.0\% 100.0100.0\%$
Visual $1.9\pm 0.7$ $85.3\pm 6.1\%$ $\pm 0.0\% 100.0100.0\%$
Wither Text $1.4\pm 0.3$ $69.5\pm 19.0\%$ $\pm 0.0\% 100.0100.0\%$
Visual $2.3\pm 0.5$ $39.1\pm 12.4\%$ $82.8\pm 7.2\%$ $66.7\%$
Ender Dragon Text $5.2\pm 1.5$ $43.9\pm 24.4\%$ $98.2\pm 5.9\%$ $91.7\%$
Visual $6.1\pm 1.1$ $21.8\pm 10.3\%$ $58.3\pm 13.5\%$ $50.0\%$
3.5. Robustness Analysis
3.5.1. Robust to Modality
We evaluated the framework’s robustness to different modalities by replacing the text-based observation with a multi-modal approach in boss combat tasks. While the inclusion of visual language models (VLMs) introduced increased response latency and slightly reduced observation accuracy, the agents maintained reasonable performance levels. As summarized in Table 4, our framework achieves strong performance across all three challenging task scenarios using either modalities.
3.5.2. Robust to Scale
To validate the effectiveness of our approach under large-scale scenarios, we conducted additional experiments involving teams of 5-50 agents. These experiments focused on measuring the impact of agent quantity (where the number of agents equaled the number of latest chat entries read per LLM inference) on token counts and inference time.
Figure 5. Inference time and input/output tokens per single-agent planning step of MAS at different scales.
In Fig. 5, it can be observed that the inference time of the LLM tends to stabilize instead of increasing continuously as the number of agents increases. Additionally, the total token cost shows an approximately linear growth with the number of agents, which is both acceptable and manageable. Since each agent is equipped with parallelized planning threads and acting threads independently, the actions of each agent remain unaffected by the total number of agents in the system. Further analysis can be found in the Appendix.
The input tokens increases linearly with the number of agents, as more chat entries are processed during each planning step. However, the output tokens remain stable due to the constraints imposed by the prompt format, which limits the length of generated responses. Overall, the system’s total latency does not increase significantly with the number of agents, and the total cost (tokens) grows approximately linearly with the number of agents, which is acceptable and controllable.
4. Related Works
@@
LLM-based Minecraft Agents. The development of LLM-based AI agents in Minecraft has evolved through several key approaches: Voyager Wang et al. (2023) established the first LLM-based agent with automatic skill discovery using GPT-4 Achiam et al. (2023), built on the open-source library Mineflayer PrismarineJS (2023). Subsequent studies enhanced agents via specialized memory mechanisms Park et al. (2024); Li et al. (2024a), specialized LLM fine-tuning Feng et al. (2023); Zhao et al. (2024); Liu et al. (2024), task decomposition and causal graph learning Yuan et al. (2023); Zhu et al. (2023); Yu and Lu (2024), and combination with reinforcement learning Li et al. (2024b, 2023). Additionally, multi-modal information perception and processing were explored Zheng et al. (2023); Cai et al. (2024), along with other novel techniques Li et al. (2025); Wang et al. (2024); Zhou et al. (2024). The development of benchmarks for general capabilities progressed with MineDojo Fan et al. (2022) and MCU Lin et al. (2023), while specific agent capabilities were assessed through additional benchmarks Qin et al. (2024); Dong et al. (2024); Wei et al. (2025).
@@
LLM-based Multi-Agent Systems. Recent research has focused on several core areas Mou et al. (2024); Guo et al. (2024): infrastructure frameworks for efficient agent coordination Hong et al. (2023); Chen et al. (2024c); Zhang et al. (2024a, b, 2025a), which introduce novel paradigms for task management and team collaboration; efforts aimed at improving MAS collaboration efficiency through optimizations at communication or routing Yue et al. (2025); Zhang et al. (2025b); Chen et al. (2025); benchmark development to evaluate multi-agent performance in dynamic environments Chen et al. (2024b); Dong et al. (2024), which has created robust testing environments to assess the generalization and efficiency of LLM-powered agents; large-scale social simulations AL et al. (2024); Park et al. (2023); Yang et al. (2024), which explore how multi-agent systems model complex societal behaviors; and domain-specific applications Wu et al. (2024); D’Arcy et al. (2024); Chen et al. (2024a) demonstrating the effectiveness of LLM-based agents in specific scenario simulation. In contrast to prior MAS research that primarily focuses on coordination paradigms or communication strategies, our work targets continuous real-time interaction in non-paused dynamic environments.
5. Conclusion
@@
We propose a novel parallelized planning-acting multi-agent framework that significantly enhances the responsiveness and adaptability of LLM-based MAS in dynamic environments like Minecraft. Our framework’s dual-thread architecture with interruptible execution mechanism enables real-time interaction and continuous adaptation, overcoming the limitations of traditional serialized execution paradigms. The comprehensive skill library and recursive task decomposition mechanism further improve efficiency and coordination as an engineering contribution. Experiments on our challenging benchmark tasks validate the effectiveness of our framework in diverse Minecraft scenarios.
{acks}
This work is supported in part by the Hangzhou Joint Funds of the Zhejiang Provincial Natural Science Foundation of China under Grant No. LHZSD24F020001, in part by the Zhejiang Province High-Level Talents Special Support Program “Leading Talent of Technological Innovation of Ten-Thousands Talents Program” under Grant No. 2022R52046, in part by the Fundamental Research Funds for the Central Universities under Grant No. 2021FZZX001-23, and in part by the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.
@@ Appendix A Centralized Memory System
The memory of the multi-agent system is stored in JSON files, and agents read and modify their contents by reading and writing to these files.
The observations and states of the agent are updated in a polling manner per second and saved in the following form, constantly overwritten.
Observation Records { time”: day”, health”: 12.8, hunger”: 17, position”: ”(10, 65, -8)”, equipment”: [ null, null, null, null, bow”, null ], inventory”: { arrow”: 187, golden_apple”: 4, diamond_pickaxe”: 1, diamond_sword”: 1, diamond_helmet”: 1, diamond_chestplate”: 1, diamond_leggings”: 1, diamond_boots”: 1, bow”: 1 }, nearbyBlocks”: [ end_stone”, bedrock ], nearbyEntities”: enderman, ender_dragon”, nearbyPlayers”: [ Steve”, Brine”, Jeb”, Notch ] }
The team chat logs are constantly updated by active communication and continue to grow in real time, but only the most recent information is taken as prompts.
@@
Chat Logs (Active Communication) { Day 1 - Morning”: [ SystemInfo: Number of End Crystals Remaining: 10”, SystemInfo: Boss Health: 200”, SystemInfo: Boss Health: 200”, Notch: Lets prioritize taking down the end_crystals to prevent the Ender Dragon from healing. I will start attacking the nearest end_crystal with my bow. Everyone, cover me and be ready to assist if needed.”, Steve: Im ready with my bow and arrows. Lets focus on taking down the end_crystals first to prevent the Ender Dragon from healing. Ill start targeting the nearest end_crystal. Stay alert for any incoming endermen but avoid engaging them unless necessary.”, SystemInfo: Boss Health: 200”, Brine: Brine: Lets prioritize destroying the end_crystals to prevent the Ender Dragon from healing. Notch and I will start targeting the nearest crystal. Steve, Alex, and Jeb, please spread out and assist in taking down other crystals. Stay alert for any Endermen but avoid engaging them unless necessary. Report if you need help or spot any crystals.”, Jeb: Im ready with my bow and arrows. Lets prioritize the end_crystal on the obsidian pillars to prevent the Ender Dragon from healing. Ill start targeting the nearest crystal. Stay alert for any incoming endermen but avoid engaging unless necessary.” ] }
The team chat logs are also continuously updated in real time by passive communication, but only the latest observations of each agent are provided to the LLM as prompts.
@@
Chat Logs (Passive Communication) { obs_entries”: [ Notch: Team, Im fully equipped with diamond armor, a bow, and 200 arrows. I see an end_crystal and an ender_dragon nearby, along with some endermen. Brine, Jeb, Steve, and Alex are also close. Lets focus on taking down the end_crystal first to prevent the Ender Dragon from healing. Everyone, prepare your bows and stay alert for any incoming endermen. Ready when you are!”, Brine: Team, Im currently well-equipped with a bow, 200 arrows, and diamond armor. The ender_dragon is nearby but hasnt attacked yet. I see Steve, Notch, Alex, and Jeb around. Lets focus on taking out the end_crystals first. I suggest we spread out to cover more ground and target the crystals from a distance. Stay alert for any Endermen, but avoid engaging them unless necessary. Report if you spot any crystals or need assistance.”, Jeb: All end_crystals have been destroyed. Focusing on the Ender Dragon now. Everyone, keep your distance and use bows to attack from a safe range. Stay alert for any incoming endermen but prioritize the Ender Dragon. Lets finish this!”, ] }
The team action logs record the historical actions and reasoning processes of each agent, but only an agent’s own current action is used as LLM prompts.
@@
In addition to this, agents are also provided with identity information such as their name, team affiliation, and teammates’ names, to simulate a human-like multi-agent collaborative scenario.
Appendix B Comprehensive Skill Library